# Structure-preserving approximation of the non-isothermal Cahn-Hilliard system

### v1

@@ 5.1 Existence of positive solutions
for all $y_{h}=(\psi_{h},\xi_{h},\omega_{h})\in\mathcal{V}_{h}\times\mathcal{V}_{h}\times\mathcal{V}_{h}$. By construction it holds $\mathbf{G}(\cdot,1)=\mathbf{g}(\cdot)=0$, hence condition $(i)$ of Theorem 1 is fulfilled.
Continuity of the map and condition (ii)
In this part we show that for every $\alpha\in[0,1]$, if a solution $\mathbf{x}\in V$ of $\mathbf{G}(\mathbf{x},\alpha)=0$ exists, then $\mathbf{x}\in W.$ To provide the required bounds of $\|\mathbf{x}\|,\leavevmode\nobreak\ \mathbf{x}=(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\vartheta})$, we follow the lines of the proof of Theorem 6 and find that for every $\alpha\in[0,1]$ we have
$\langle e(\phi_{h},\vartheta_{h})-e(\phi_{h}^{n},\vartheta_{h}^{n}),1\rangle$ $=\tau\mathcal{D}_{num}^{n+1}\leq 0,$
$\langle s(\phi_{h},\vartheta_{h})-s(\phi_{h}^{n},\vartheta_{h}^{n}),1\rangle$ $=\tau\alpha\mathcal{D}_{h}^{n+1}\geq 0.$
Combination yields
$\langle e(\phi_{h},\vartheta_{h})-c_{f}s(\phi_{h},\vartheta_{h}),1\rangle$ $-\tau\mathcal{D}_{num}^{n+1}+c_{f}\tau\alpha\mathcal{D}_{h}^{n+1}$
$=\langle e(\phi_{h}^{n},\vartheta_{h}^{n})-c_{f}s(\phi_{h}^{n},\vartheta_{h}^{n}),1\rangle\leq C^{\prime}.$
First, since $(\phi_{h},\mu_{h},\vartheta_{h})\in V$ we find that $\vartheta_{h}>0$ and thus there exists $c_{1},c_{2}$ such that $c_{1}\leq\vartheta_{h}\leq c_{2}.$ Note that the lower and upper bounds may depend on $\alpha$. To obtain independent bounds we take the minimum/maximum with respect to $\alpha.$ Furthermore using $\psi_{h}=1$ we derive conservation of mass and hence control on the mean value of $\phi_{h}.$ Using (A5) and Poincare inequality for $\phi_{h}$ we obtain
$\|\phi_{h}\|_{H^{1}}^{2}+\|\vartheta_{h}\|_{L^{1}(\Omega)}\leq C^{*}C^{\prime},$ (20)
where $C^{\prime}$ only depends on the norms of known functions $(\phi_{h}^{n},\vartheta_{h}^{n}).$ It remains to obtain an estimate on $\mu_{h}$:
$\|\mu_{h}\|_{L^{2}}^{2}$ $=\gamma\langle\vartheta_{h}\nabla\phi_{h},\nabla\mu_{h}\rangle+\gamma\langle\nabla\phi_{h}^{n}\nabla\vartheta_{h},\mu_{h}\rangle+\langle f_{\phi}(\phi_{h},\phi_{h}^{n},\vartheta_{h}),\mu_{h}\rangle$
$\leq\gamma c_{2}\|\nabla\phi_{h}\|_{L^{2}}\|\nabla\mu_{h}\|_{L^{2}}+\gamma\|\nabla\phi_{h}^{n}\|_{L^{2}}\|\nabla\vartheta_{h}\|_{L^{2}}\|\mu_{h}\|_{L^{\infty}}$
$\;+\|f_{\phi}(\phi_{h}^{n+1},\phi_{h}^{n},\vartheta_{h}^{n+1})\|_{L^{1}}\|\mu_{h}\|_{L^{\infty}}$
$=(a)+(b)+(c).$
Using the inverse inequality (13) we find that there exists a constant $C=C(h^{-1})$ such that
$(a)$ $\leq C(h^{-1})\|\nabla\phi_{h}\|_{L^{2}}\|\mu_{h}\|_{L^{2}},$
$(b)$ $\leq C(h^{-1})\|\nabla\phi_{h}^{n}\|_{L^{2}}\|\vartheta_{h}\|_{L^{1}}\|\mu_{h}\|_{L^{2}},$
$(c)$ $\leq C(h^{-1})\|f_{\phi}(\phi_{h}^{n+1},\phi_{h}^{n},\vartheta_{h}^{n+1})\|_{L^{1}}\|\mu_{h}\|_{L^{2}}.$
For the last term we use the lower and upper bounds for $\vartheta_{h}$ and (A6) to estimate
$(c)\leq C(\|\phi_{h}\|_{H^{1}}^{p_{1}}+\|\phi_{h}^{n}\|_{H^{1}}^{p_{2}}+\|\vartheta_{h}\|_{L^{1}}^{k}+1)\|\mu_{h}\|_{L^{2}}.$
Using (a), (b), (c), the lower/upper bounds for $\vartheta$ and the bounds (20) we find
$\|\mu_{h}\|_{L^{2}}\leq\bar{C}.$ (21)
Hence, we choose the constants $C_{\phi,\mu},\varepsilon,C_{\vartheta}$ in (16) as
[…]
Condition (iii): Topological degree of $\mathbf{G}(\cdot,0)$:
@@
In this part we show that the problem $\mathbf{G}(\mathbf{x},0)=0$ has a solution $\mathbf{x}\in W$ and that the topological degree of $\mathbf{G}(\cdot,0)$ is non-zero. We start by reformulating $\mathbf{G}(\mathbf{x},0)$, i.e. (17)–(19) with $\alpha=0$, which reads
@@
$\langle\mu_{h},\xi_{h}\rangle-\gamma\langle\nabla\phi_{h},\vartheta_{h}\nabla\xi_{h}\rangle-\gamma\langle\nabla\phi_{h}^{n},\xi_{h}\nabla\vartheta_{h}\rangle-\langle\partial_{\phi}f_{h},\xi_{h}\rangle$ $=0$ (23)
@@
Next, we will show that the topological degree is non-zero. To this end we compute the Jacobian of $\mathbf{G}(\cdot,0)$
$\mathbf{J}(\mathbf{G}(\cdot,0))=\begin{pmatrix}\mathbf{A}&\mathbf{0}&\mathbf{0}\\ \mathbf{H}^{\phi}&\mathbf{A}&\mathbf{H}^{\theta}\\ \mathbf{S}^{\phi}&\mathbf{0}&\mathbf{S}^{\vartheta}\end{pmatrix}$
@@
$\mathbf{A}_{ij}$ $=\langle\lambda_{j},\lambda_{i}\rangle,\quad\mathbf{H}_{i,j}^{\phi}=\gamma\langle\nabla\lambda_{j},\vartheta_{h}\nabla\lambda_{i}\rangle+\langle\partial_{\phi_{h}}f_{\phi}(\phi_{h},\phi_{h}^{n},\vartheta_{h})\lambda_{j},\lambda_{i}\rangle,$
@@
$\mathbf{H}_{i,j}^{\theta}=$ $\gamma\langle\nabla\phi_{h},\lambda_{j}\nabla\lambda_{i}\rangle+\gamma\langle\nabla\phi_{h}^{n},\lambda_{i}\nabla\lambda_{j}\rangle+\langle\partial_{\vartheta_{h}}f_{\phi}(\phi_{h},\phi_{h}^{n},\vartheta_{h})\lambda_{j},\lambda_{i}\rangle,$
$\mathbf{S}^{\phi}_{i,j}$ $=\langle\partial_{\phi}s(\phi_{h},\vartheta_{h})\lambda_{j},\lambda_{i}\rangle,\quad\mathbf{S}^{\vartheta}_{i,j}=\langle\partial_{\vartheta}s(\phi_{h},\vartheta_{h})\lambda_{j},\lambda_{i}\rangle.$
@@
Here the functions $\lambda_{i}$ are the finite element basis functions. The topological degree involves the signum of the determinant of the Jacobian. We can compute that $\det(\mathbf{J}(\mathbf{G}(\cdot,0)))=\det(\mathbf{A})^{2}\det(\mathbf{S}^{\vartheta}).$ The mass matrix $\mathbf{A}$ is positive definite, hence $\det(\mathbf{A})^{2}>0$. Since $\partial_{\vartheta}s=-\partial_{\vartheta\vartheta}F$, we find using strict concavity assumptions of $f$, i.e (A5), that $-\partial_{\vartheta\vartheta}F>0$ and therefore $\mathrm{sgn}(\det(\mathbf{J}(\mathbf{G}(\mathbf{x},0))))=1$. The topological degree is given by
$\mathrm{deg}(\mathbf{G}(\cdot,0)):=\sum_{\mathbf{x}\in\mathbf{G}^{-1}(0)}\mathrm{sgn}(\det(\mathbf{J}(\mathbf{G}(\mathbf{x},0))))=\sum_{\mathbf{x}\in\mathbf{G}^{-1}(0)}1.$
@@
Hence, the topological degree is non-zero, if at least one $\mathbf{x}\in\mathbf{G}^{-1}(0)$ exists. We have already shown that system (22)–(24) admits at least one solution. Hence, the topological degree is non-zero.
@@
Finally, we can apply Theorem 1 for one time step of the problem $\mathbf{G}(\mathbf{x},\alpha)=0$ and find that at least one solution $\mathbf{x}\in W$ exists. The existence of at least one discrete solution $x_{h}=(\phi_{h},\mu_{h},\vartheta_{h})$ can be done in an inductive way. ∎
@@
We should note that Theorem 7 gives the existence of a bounded solution with strictly positive temperature. The bounds depend on $h,\tau$ due to the use of inverse inequalities. Note that the quasi-uniformity assumptions is only for simplicity and can be omitted, if local mesh sizes are considered.

### v2

@@ 5.1 Existence of positive solutions
By construction it holds $\mathbf{G}(\cdot,1)=\mathbf{g}(\cdot)=0$, hence condition $(i)$ of Theorem 1 is fulfilled.
Condition (ii): Topological degree of $\mathbf{G}(\cdot,0)$:
@@
In this part we show that the problem $\mathbf{G}(\mathbf{x},0):=\mathbf{G}_{0}(\mathbf{x})=0$ has a solution $\mathbf{x}\in W$ and that the topological degree of $\mathbf{G}_{0}$ is non-zero. We start by formulating $\mathbf{G}_{0}(\mathbf{x})=0$, i.e. (21)–(23) with $\alpha=0$, which reads
@@
$\langle\mu_{h},\xi_{h}\rangle-\gamma\langle\nabla\phi_{h},\vartheta_{h}\nabla\xi_{h}\rangle-\gamma\langle\nabla\phi_{h},\xi_{h}\nabla\vartheta_{h}\rangle-\langle f_{\phi}(\phi_{h},\phi_{h}^{n},\vartheta_{h}),\xi_{h}\rangle$ $=0$ (25)
@@
$\langle\partial_{\vartheta}s(\phi_{h}^{n},\zeta)(\vartheta_{h}-\vartheta_{h}^{n}),\omega_{h}\rangle=0,\qquad\zeta\in[\vartheta_{h}^{n},\vartheta_{h}].$
Recalling that $\partial_{\vartheta}s=-\partial_{\vartheta\vartheta}F$ and strict concavity of $F$, i.e. (A5), we get that $\vartheta_{h}=\vartheta_{h}^{n}$. Then $\mu_{h}$ can be computed from (25). It remains to show that $(\phi_{h},\mu_{h},\vartheta_{h})\in W$. In fact it only remains to show the bound on $\|\mu_{h}\|$. Hence, we estimate as follows:
$\|\mu_{h}\|_{L^{2}}^{2}$ $=\gamma\langle\vartheta_{h}\nabla\phi_{h},\nabla\mu_{h}\rangle+\gamma\langle\nabla\phi_{h}\nabla\vartheta_{h},\mu_{h}\rangle+\langle f_{\phi}(\phi_{h},\phi_{h},\vartheta_{h}),\mu_{h}\rangle$
$\leq\gamma c_{2}\|\nabla\phi_{h}\|_{L^{2}}\|\nabla\mu_{h}\|_{L^{2}}+\gamma\|\nabla\phi_{h}\|_{L^{2}}\|\nabla\vartheta_{h}\|_{L^{2}}\|\mu_{h}\|_{L^{\infty}}$
$\;+\|f_{\phi}(\phi_{h},\phi_{h},\vartheta_{h})\|_{L^{1}}\|\mu_{h}\|_{L^{\infty}}$
$=(a)+(b)+(c).$
Using the inverse inequality (17) we find that there exists constants $C_{3},C_{4},C_{5}>0$ such that
$(a)$ $\leq C_{3}h^{-1}\|\nabla\phi_{h}\|_{L^{2}}\|\mu_{h}\|_{L^{2}},$
$(b)$ $\leq C_{4}h^{-(d+1)}\|\nabla\phi_{h}\|_{L^{2}}\|\vartheta_{h}\|_{L^{1}}\|\mu_{h}\|_{L^{2}},$
$(c)$ $\leq C_{5}h^{-d/2}\|f_{\phi}(\phi_{h},\phi_{h},\vartheta_{h})\|_{L^{1}}\|\mu_{h}\|_{L^{2}}.$
For (c) use assumption (A6) as well as the Hölder and Sobolev inequalities to estimate
$\|f_{\phi}(\phi_{h},\phi_{h},\vartheta_{h})\|_{L^{1}}$ $\leq C_{f}(\bar{\vartheta}^{n}+1)(2\|\phi_{h}\|_{H^{1}}^{6}+|\Omega|)\|\mu_{h}\|_{L^{2}}.$
Since $\phi_{h}$ and $\vartheta_{h}$ are bounded we find
$\|\mu_{h}\|_{L^{2}}\leq C_{\mu}.$ (27)
Next, we will show that the topological degree is non-zero. To this end we compute the Jacobian matrix of $\mathbf{G}_{0}$
$J_{\mathbf{G}_{0}}(\mathbf{x})=\begin{pmatrix}\mathbf{A}&\mathbf{0}&\mathbf{0}\\ \mathbf{H}^{\phi}&\mathbf{A}&\mathbf{H}^{\theta}\\ \mathbf{S}^{\phi}&\mathbf{0}&\mathbf{S}^{\vartheta}\end{pmatrix}$
@@
$\mathbf{A}_{ij}=\langle\lambda_{j},\lambda_{i}\rangle\quad\mathbf{S}^{\phi}_{i,j}=\langle\partial_{\phi}s(\phi_{h},\vartheta_{h})\lambda_{j},\lambda_{i}\rangle,\quad\mathbf{S}^{\vartheta}_{i,j}=\langle\partial_{\vartheta}s(\phi_{h},\vartheta_{h})\lambda_{j},\lambda_{i}\rangle,$
@@
$\mathbf{H}_{i,j}^{\phi}=\gamma\langle\nabla\lambda_{j},\vartheta_{h}\nabla\lambda_{i}\rangle+\gamma\langle\nabla\lambda_{j},\lambda_{i}\nabla\vartheta_{h}\rangle+\langle\partial_{\phi_{h}}f_{\phi}(\phi_{h},\phi_{h}^{n},\vartheta_{h})\lambda_{j},\lambda_{i}\rangle,$
@@
$\mathbf{H}_{i,j}^{\theta}=\gamma\langle\nabla\phi_{h},\lambda_{j}\nabla\lambda_{i}\rangle+\gamma\langle\nabla\phi_{h},\lambda_{i}\nabla\lambda_{j}\rangle+\langle\partial_{\vartheta_{h}}f_{\phi}(\phi_{h},\phi_{h}^{n},\vartheta_{h})\lambda_{j},\lambda_{i}\rangle.$
@@
Here the functions $\lambda_{i}$ are the finite element basis functions. The topological degree involves the signum of the determinant of the Jacobian matrix. We can compute that $\det(J_{\mathbf{G}_{0}}(\mathbf{x}))=\det(\mathbf{A})^{2}\det(\mathbf{S}^{\vartheta}).$ The mass matrix $\mathbf{A}$ is positive definite, hence $\det(\mathbf{A})^{2}>0$. Since $\partial_{\vartheta}s=-\partial_{\vartheta\vartheta}F$, we find using strict concavity assumptions of $F$, i.e (A5), that $\partial_{\vartheta\vartheta}F>0$ and therefore $\mathrm{sgn}(\det(J_{\mathbf{G}_{0}}(\mathbf{x})))=1$. The topological degree is then given by
$\mathrm{deg}(\mathbf{G}_{0},W,\mathbf{0}):=\sum_{\mathbf{x}\in\mathbf{G}_{0}^{-1}(\mathbf{0})}\mathrm{sgn}\left(\det\left(J_{\mathbf{G}_{0}}(\mathbf{x})\right)\right)=\sum_{\mathbf{x}\in\mathbf{G}_{0}^{-1}(\mathbf{0})}1.$
@@
Hence, the topological degree is non-zero, if at least one $\mathbf{x}\in\mathbf{G}_{0}^{-1}(0)$ exists. We have already shown that system (24)–(26) admits at least one solution. Consequently, the topological degree is non-zero.
Moreover, in this case the solution satisfies the assumptions of the implicit function theorem and hence there exists $\delta>0$ such that for all $\alpha\in[0,\delta)$ we have that $\mathbf{x}=z(\alpha)$, for a continuously differentiable function $z$. Due to the continuity of $\vartheta$ with respect to $\alpha$, the minimum of $\vartheta_{h}$ remains strictly positive for all $\alpha\in[0,\tfrac{\delta}{2}]$. We denote the minimal value of $\vartheta_{h}$ for $\alpha\in[0,\tfrac{\delta}{2}]$ by $C_{\delta}>0$.
Continuity of the map and condition (iii)
In this part we show that for every $\alpha\in[0,1]$, if a solution $\mathbf{x}\in V$ of $\mathbf{G}(\mathbf{x},\alpha)=0$ exists, then $\mathbf{x}\in W.$ To provide the required bounds of $\|\mathbf{x}\|,\penalty 10000\ \mathbf{x}=(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\vartheta})$, we follow the lines of the proof of Theorem 6 and find that for every $\alpha\in[0,1]$ we have
$\langle\phi_{h}-\phi_{h}^{n},1\rangle$ $=0$ (28)
$\langle e(\phi_{h},\vartheta_{h})-e(\phi_{h}^{n},\vartheta_{h}^{n}),1\rangle$ $=\tau\mathcal{D}_{num}\leq 0,$ (29)
$\langle s(\phi_{h},\vartheta_{h})-s(\phi_{h}^{n},\vartheta_{h}^{n}),1\rangle$ $=\tau\alpha\mathcal{D}_{h}\geq 0.$ (30)
Combination of the latter two yields
$\langle e(\phi_{h},\vartheta_{h})-c_{f}s(\phi_{h},\vartheta_{h}),1\rangle$ $-\tau\mathcal{D}_{num}+c_{f}\tau\alpha\mathcal{D}_{h}$
$=\langle e(\phi_{h}^{n},\vartheta_{h}^{n})-c_{f}s(\phi_{h}^{n},\vartheta_{h}^{n}),1\rangle\leq C_{e,s}$ (31)
[…]
$\langle e(\phi_{h},\vartheta_{h})-c_{f}s(\phi_{h},\vartheta_{h}),1\rangle$ $=\langle b(\vartheta_{h}-c_{f}-c_{f}\log(\vartheta_{h}))$
$\quad+f_{2}(\phi_{h})-c_{f}f_{1}(\phi_{h})+c_{f}\tfrac{\gamma}{2}|\nabla\phi_{h}|^{2},1\rangle.$ (32)
Because $b(\vartheta_{h}-c_{f}-c_{f}\log(\vartheta_{h}))$ is bounded from below by $bc_{f}\log(c_{f})$ for any $\vartheta_{h}$ and $f_{2}(\phi_{h})-c_{f}f_{1}(\phi_{h})>-c$, we obtain
$\|\nabla\phi_{h}\|_{L^{2}}^{2}$ $\leq 2\frac{C_{e,s}+|\Omega|\left(bc_{f}\log(c_{f})+c\right)}{c_{f}\gamma}.$
Applying (28) and the Poincaré inequality for $\phi_{h}$ we have
$\|\phi_{h}\|_{H^{1}}^{2}\leq C_{\phi}.$ (33)
To derive an upper bound for $\vartheta_{h}$ we observe that the internal energy has the form $e=\vartheta_{h}+f_{2}(\phi_{h})$ with $f_{2}(\phi_{h})>-c$. Using the positivity of $\vartheta_{h}$ and (29) we find that $\|\vartheta_{h}\|_{L^{1}}\leq C_{\|\vartheta\|}$. Considering again (32) yields
$\int_{\Omega}-\log(\vartheta_{h})\leq\tfrac{c+C_{e,s}}{bc_{f}}+1=:C_{\log\vartheta}.$ (34)
These bounds together with (33) are sufficient to compute the uniform bound for $\mu_{h}$ in the same spirit of (ii), i.e. $\|\mu_{h}\|_{L^{2}}\leq C_{\mu}$, for $C_{\mu}>0$. Hence, it remains to show the uniform lower bound of $\vartheta_{h}$.
Recalling (31) and the uniform positive definiteness assumption on the Onsager matrix $\mathbf{L}(\phi_{h},\nabla\phi_{h},\vartheta_{h})$, i.e. (A3), we find
$c_{f}\tau\alpha\lambda_{0}\left(\left\|\nabla\frac{1}{\vartheta_{h}}\right\|_{L^{2}}^{2}+\left\|\nabla\frac{\mu_{h}}{\vartheta_{h}}\right\|_{L^{2}}^{2}\right)\leq c_{f}\tau\alpha\mathcal{D}_{h}\leq C_{e,s}.$
We consider two different situations. First, we consider an element $K$ on which $\nabla\vartheta_{h}$ is non-zero and estimate
$\frac{C_{e,s}}{c_{f}\tau\alpha\lambda_{0}}\geq\int_{K}\left|\nabla\frac{1}{\vartheta_{h}}\right|^{2}=|\nabla\vartheta_{h}|^{2}\int_{K}\frac{1}{\vartheta_{h}^{4}}\geq C_{\nabla\vartheta}\|\vartheta_{h}^{-1}\|_{L^{4}(K)}^{4}.$
As the gradient is by assumption non-zero we can restrict the integration domain from $K$ to $\tilde{K}:=\{x\in K:\vartheta_{h}(x)\in[\vartheta_{\min},2\vartheta_{\min}]\}$, where $\vartheta_{\min}$ is the minimum on $K$ and $|\tilde{K}|>0$. By construction we obtain
$\vartheta_{\min}\geq\left(\frac{C_{\nabla\vartheta}c_{f}\tau\alpha\lambda_{0}|\tilde{K}|}{8C_{e,s}}\right)^{1/4}\geq\left(\frac{C_{\nabla\vartheta}c_{f}\tau\delta\lambda_{0}|\tilde{K}|}{16C_{e,s}}\right)^{1/4}:=C_{\text{grad}}.$
Second, we consider the case when $\vartheta_{h}=\vartheta_{\text{const}}$ is constant over one element $K$, i.e. $\nabla\vartheta_{h}|_{K}=0$. Then either $\vartheta_{h}=\vartheta_{\text{const}}$ over $\Omega$ or there exists an element $\bar{K}$ on which $\nabla\vartheta_{h}\neq 0$ and $\vartheta_{h}(x_{j})=\vartheta_{\text{const}}$ at least in one nodal point $x_{j}$. If $\vartheta_{h}$ is constant in $\Omega$, we obtain using (34) that
$\vartheta_{h}\geq\exp\left(-\frac{C_{\log\vartheta}}{|\Omega|}\right).$
In the other case, i.e. $\nabla\vartheta_{h}\neq 0$ on some element, we use the first result. Hence, we have derived the uniform lower bound
$\vartheta_{h}\geq\min\left(\exp\left(-\frac{C_{\log\vartheta}}{|\Omega|}\right),C_{\text{grad}}\right)=:C_{\text{min}}$
for $\alpha\in[\tfrac{\delta}{2},1]$. To obtain a uniform bound for $\alpha\in[0,1]$ we take the minimum of $C_{\text{min}}$ and $C_{\delta}$. We choose the constants $C_{\phi,\mu},C_{\vartheta},\varepsilon$ in (20) in the following way
$C_{\phi,\mu}=2C_{\|\cdot\|}\max\{\sqrt{C_{\phi}},C_{\mu}\},\;C_{\vartheta}=2\max\{C_{\|\cdot\|_{\infty}}C_{\|\vartheta\|},\bar{\vartheta}^{n}\},\;\varepsilon=\frac{1}{2}\min\{C_{\text{min}},C_{\delta}\}.$
Because $C_{\|\vartheta\|}$ is the bound in the discrete $L^{1}$ norm, we multiply by the norm equivalence constant $C_{\|\cdot\|_{\infty}}$. The same applies to the $L^{2}$ bound $C_{\phi,\mu}$ with the constant $C_{\|\cdot\|}$.
@@
Finally, we can apply Theorem 1 for one time step of the problem $\mathbf{G}(\mathbf{x},\alpha)=0$ and find that at least one solution $\mathbf{x}\in W$ for $\mathbf{g}(\mathbf{x})=0$ exists. The existence of at least one discrete solution $(\phi_{h},\mu_{h},\vartheta_{h})$ can be done in an inductive way. ∎
@@
  1. 1.
    We should note that Theorem 8 gives the existence of a bounded solution with strictly positive temperature. The bounds depend on $h,\tau$ due to the use of inverse inequalities. Note that the quasi-uniformity assumptions is only for simplicity and can be omitted, if local mesh sizes are considered. Furthermore, we expect the result to be extendable to a larger class of Helmholtz free energies $f$.
  2. 2.
    Uniqueness of solution for small $\tau$ is a consequence of the implicit function theorem. However, for larger $\tau$ multiple solution may exist and can affect the convergence of the Newton method.